OpenStudy (anonymous):

Find the solution of the exponential equation: 2^(2x+6) = 3^(x-39) in terms of logarithms, or correct to four decimal places.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

be prepared to do a fair amount of algebra here ready?

OpenStudy (anonymous):

yass

OpenStudy (misty1212):

ok we start with \[(2x+6)\ln(2)=(x-39)\ln(3)\] then solve for \(x\) with the understanding that \(\ln(2)\) and \(\ln(3)\) are just some numbers (constants)

OpenStudy (misty1212):

this is where the algebra comes in multiply out \[2\ln(2)x+6\ln(2)=\ln(3)x-39\ln(3)\] good so far?

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

k now we have to put all the x terms on one side of the equal sign, all the numbers on the other

OpenStudy (misty1212):

\[2\ln(2)x-\ln(3)x=-6\ln(2)-39\ln(3)\]

OpenStudy (misty1212):

next factor the \(x\) out of the left hand side so we know what to divide by \[(2\ln(2)-\ln(3))x=-6\ln(2)-39\ln(3)\]

OpenStudy (misty1212):

and finally divide \[x=\frac{-6\ln(2)-39\ln(3)}{2\ln(2)-\ln(3)}\]

OpenStudy (misty1212):

now use a calculator, since you have to get it in decimal form

OpenStudy (anonymous):

ok thnx

OpenStudy (misty1212):

on the other hand, since we have to use a calculator anyways, we might as well just solve it with a calculator

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=2%5E%282x%2B6%29+%3D+3%5E%28x-39%29 click on "approximate form" and you will see the answer

OpenStudy (anonymous):

i already did it and it was the right answer, thank you so much